With the info provided make a calculation of the direct cost, direct labor and determine the base cost of raw material and the labor base cost according to the following lists.
The company KIA motors SA de CV presents the following materials:
1. #12 steel sheets, $ 1500 each.
2. High quality industrial iron, $ 45,000 per ton
3. Industrial rubber # 23 at a cost of $ 13,500 per ton
4. Electronic sensors package for a unit at a cost of $ 4,500.
5. Set of lights for one unit $ 12,500
6. Engine of 5 horses of 16 valves, 33,500 per unit
7. Game of internal accessories of the vehicle, $ 1,200 per unit package
8. Set of screws for a total assembly, $ 9,800 per game
9. Industrial nylon for protection of the units, $ 4,500 per ton
10. Transfer of finished units to the warehouse, $ 250 per unit
Labour expenses
1. 10 specialized technicians, each one receives a salary of $ 4,500
2.15 auxiliary technicians, each one receives $ 750 per day.
3. A plant supervisor with a monthly salary of $ 35,000
4. 3 auxiliary supervisors with a monthly salary of $ 15,000
5. The technicians generate a social security tax of 26% monthly on salaries.
6. A 5% monthly tax is paid to Health insurance.
7. A monthly payroll tax is paid on the technicians.

8. Finally it takes some unforeseen extra hours for auxiliary technicians of 9 hours per week.

1.Determine the base cost of the car per unit if for the case of the materials 10 cars are produced with the materials of point 2, 3 and 9 and for the case of point 1,4 steel sheets are used, in all the others a set of units.
2. Determine the base cost of labor and consider that for the construction of a car they do it in 3 hours a day.
3. Determine the total production cost if 300 cars are produced per month.
Answer #1

ได้ 145 000 Что JADI S000 1350 13800 33S00 1 O0 I 14030

2. Base cost of Labor calculations

Given that 300 cars are produced per month. Also that 1 car requires 3 hours of labor.

So, Labor hours per month = 300 cars x 3 hours per car = 900 hours per type of labor.

We'll first compute the labor cost per car for each labor separately.

Specialized Technicians

10 technicians
monthly salary = $ 4500 per technician
Social Security tax = 26% on Salary = 26% of $4500 = $1170
Health Insurance = 5% of Salary = 5% of $4500 = $225
Total costs of 1 Specialized technician per month = $4500 + $1170 + $225 = $5895
Total costs towards 10 Technicians = $58,950
900 hours of work is done per month.
So, Labor cost of Specialized Technicians per hour = $58,950 / 900 hours per month = $65.5 per hour
Auxiliary Technicians

In the absence of information of working hours per day and working days per month, we will assume that the auxiliary technicians also work for the same time per day as Specialized technicians work( which is 9 hours per day)
So, each auxiliary technician works 9 hours a day
15 technicians (as the pay is per technician per day, number of technicians become irrelevant here, as we will make calculations in terms of number of technician days)
Daily pay = $ 750 per technician
Work required per month = (3 hours x 300 cars) + (9 unforeseen hours x 4 weeks) = 900 + 36 = 936 hours
Number of work days required per month = 936 hours / 9 hours per day = 104 technician days
Monthly pay to Auxiliary Technicians = 104 technician days x $750 per day = $78,000
Social Security tax = 26% on $78000 = $20,280
Health Insurance = 5% of $78,000 = $3900
Total costs towards Auxiliary technicians per month = $78000 + $20280 + $3900 = $102180
So, Labor cost of Auxiliary Technicians per hour = $102180 / 936 hours per month = $109.17 per hour
Plant Supervisor

1 Supervisor
monthly salary = $ 35000 per month
Social Security tax = Payable only for technicians, not supervisors
Health Insurance = 5% of Salary = 5% of $35000 = $1750
Total costs towards 1 Plant Supervisor per month = $35000 + $1750 = $36750
900 hours of work is done per month.
So, Labor cost of Plant Supervisor per hour = $36750 / 900 hours per month = $40.83 per hour
Auxiliary Supervisor

3 Supervisors
monthly salary = $ 15000 per month
Social Security tax = Payable only for technicians, not supervisors
Health Insurance = 5% of Salary = 5% of $15000 = $750
Total costs towards 3 Auxiliary Supervisors per month = ($15000 + $750) x 3 = $47250
900 hours of work is done per month.
So, Labor cost of Plant Supervisor per hour = $47250 / 900 hours per month = $52.50 per hour
Base Cost of Labor per hour96.50